# Cohomology ring of non-compact abelian arrangements

### v1

@@ Abstract.
We give a Orlik-Solomon type presentation for the cohomology ring of arrangements in a non-compact abelian Lie group. The new insight consists in comparing arrangements in different abelian groups. Our work is based on the Varchenko-Gelfand ring for real hyperplane arrangements and from that we deduce the cohomology rings of all other abelian arrangements. As by-product, we obtain a new proof of the Orlik-Solomon relations and De Concini-Procesi ones.
@@ 1. Introduction
The topology of hyperplane arrangements is a classical subject whose study goes back to the sixties. It is originally motivated by the relations with braid groups [Del72] and configuration spaces [CLM76]. The study of arrangements was also motivated by combinatorics, in particular matroid theory, and by singularity theory. In the last few years, the new development in matroid theory renewed the interest in the topology of arrangements and their compactifications: June Huh proved inequalities for invariants of realizable matroids by using properties of hyperplane arrangements [Huh12], that were later generalized to all matroids [AHK18]. Recently, hyperplane arrangements were used as local model in algebraic geometry, especially in moduli problems [dCHM21, MMP23].
In the 2000s, De Concini and Procesi [DCP05] started the study of more general arrangements – called toric arrangements – motivated by the application to knapsack problem. Finally, abelian arrangements were introduced in 2016 by Bibby [Bib16].
An important focus on this subject is the description of the cohomology ring of the arrangement complement. Arnold identified relations between differential forms in the cohomology ring of braid hyperplane arrangements [Arn69] and Brieskorn described the cohomology ring of braid arrangements [Bri72]. A milestone is the famous result by Orlik and Solomon [OS80]: they presented the cohomology ring of complex hyperplane arrangements by generators and relations. Varchenko and Gelfand studied the cohomology ring of real hyperplane arrangements using Heaviside functions as generators [VG87]. They observed an analogy between the real and the complex case:
@@
Proudfoot [Pro06] relates the VG-algebra and the OS-algebra with $\mathbb{F}_{2}$-coefficients overcoming the non-commutativity of the OS algebra. Moseley [Mos17] shows an isomorphism between the VG-algebra and the cohomology of (subspace) arrangements in $(\mathbb{R}^{3})^{n}$. The Varchenko-Gelfand presentation was generalized to oriented matroids and conditional oriented matroids [GR89, Cor02, DB23, DBPW22].
A Orlik-Solomon type presentation for toric arrangements was obtained by De Concini and Procesi [DCP05] in the unimodular case and later extended to all toric arrangements [CD17, ${}^{+}$20]. The problem of understanding arrangements of subtori was approached in a sequence of articles [DCG18, DCG19, DCGP20, DCG21, MP22].
@@
We obtain a new proof of the Orlik-Solomon result [OS80] for complexified hyperplane arrangements and of the De Concini-Procesi presentation [DCP05] for unimodular toric arrangements. In all other cases our result is original.
We start by setting the notations in Section 2. In Section 3, we explicitly describe the pushforward along the map $i\colon M^{\mathbb{R}}(\mathcal{A})\to M^{a,b}(\mathcal{A})$ induced by the inclusion $\mathbb{R}\hookrightarrow\mathbb{R}^{b}\times(S^{1})^{a}$. We manipulate the Varchenko-Gelfand relations to obtain relations in the cohomology ring of central and unimodular arrangements (see Lemma 3.12). Section 4 extends the results of [LTY21] to the case of non-central arrangements (see Theorem 4.1). Then, in Section 5 we use the technique of separating coverings developed in [${}^{+}$20] to obtain the main theorem. Finally, in Section 6 we apply our theorem to the arrangements of type $A_{n}$ improving a result of [CT78] on configuration spaces in $\mathbb{R}^{b}\times(S^{1})^{a}$.
@@ 2. Definitions and notations
Let $\Lambda\cong\mathbb{Z}^{r}$ be a lattice with the choice of an orientation, so that $G^{r}$ has a natural orientation. A primitive element $\chi\in\Lambda$ corresponds to a group morphism $\chi\colon G^{\Lambda}\rightarrow G$. Let $E$ be a finite set with a total ordering. An abelian arrangement is a finite collection of connected subvarieties in $G^{\Lambda}$ defined by a finite list of functions $\chi_{i}$ and elements $g_{i}\in G$ indexed by $E$:
@@
This definition, specialized to the case $G=\mathbb{C}$, corresponds to the definition of complexified hyperplane arrangements in the central case; in the non central case we allow also complex translations.
The poset of layers (or intersection poset) of $\mathcal{A}$, $\mathcal{L}(\mathcal{A})$, is defined as the set of all connected components of intersections of elements in $\mathcal{A}$, ordered by reverse inclusion. An arrangement is called central if the intersection of all the subvarieties is non-empty, and it is called unimodular whenever all the possible intersections of elements in $\mathcal{A}$ is either connected or empty.
@@
$M(\mathcal{A})=G^{\Lambda}\setminus\bigcup_{i\in E}H_{i}.$
@@
In order to define some classes in $H^{a+b-1}(G\backslash\{e\})$ we consider the immersion of a small sphere $i\colon S^{a+b-1}\hookrightarrow G\setminus e$ centered in $e$. Let $W$ be the top dimensional class generating $H_{a+b-1}(S^{a+b-1})$. We denote by $\omega$ the class in $H^{a+b-1}(G\setminus\{e\})$ dual to $i_{*}(W)$ with the orientation given by the outer normal first rule.
@@
$i_{x}\colon\quad(S^{1})^{a}$ $\hookrightarrow G\setminus\{e\}=(\mathbb{R}^{b}\times(S^{1})^{a})\setminus(0,1,\cdots,1)$
@@
Up to homotopy, $i_{x}$ only depends on the connected component of $\mathbb{R}^{b}\setminus\{0\}$ $x$ belongs to. In the case $b=1$ we choose $x<0$, otherwise the choice of the point does not matter. Denote by $\psi^{j}$ the class in $H^{1}(G\setminus\{e\})$ dual to $(i_{x})_{*}(Y^{j})$ and by $\psi$ the class in $H^{a}(G\setminus\{e\})$ dual to $(i_{x})_{*}(Y)$.
For any $i\in E$, the map $\chi_{i}$ restricts to $(\chi_{i}-g_{i})|_{M^{a,b}(\mathcal{A})}\colon M^{a,b}(\mathcal{A})\rightarrow G\setminus\{e\}$. So we set
@@
Note that, when $b=1$, the classes $\psi_{i}$’s have degree $a=a+b-1$, the same degree of the elements $\omega_{i}$’s.
@@ 3. Central and unimodular arrangements
Consider a proper map $f\colon N\to M$ between oriented manifolds, it induces a pushforward map in cohomology $f_{*}$ obtained from the pushforward in Borel-Moore homology (for a general reference [HR96]) by composing with the Poincaré duality isomorphism.
{tikzcd}
@@
We consider the pushforward in cohomology $i_{*}\colon H^{0}(\mathbb{R}\setminus\{0\})\rightarrow H^{a+b-1}(G\setminus\{e\})$ induced by the closed inclusion $i\colon\mathbb{R}\setminus\{0\}\rightarrow G\setminus\{e\}$, $i(x)=(x,0,\cdots,0,1,\cdots,1)$. Let $w^{+}$ and $w^{-}$ be the two standard generators of $H^{0}(\mathbb{R}\setminus\{0\})$. Poincaré Duality maps $w^{+}$ and $w^{-}$ to the classes in $H^{\operatorname{BM}}_{1}(\mathbb{R}\setminus\{0\})$ represented by the infinite chains $c^{+}$ and $c^{-}$, corresponding respectively to the positive and negative semi-axes. By checking intersection numbers, we obtain $\operatorname{PD}(\omega)=j_{*}c^{+}$ and, when $b=1$, $\operatorname{PD}(\psi)=j_{*}c^{+}+j_{*}c^{-}=j_{*}1$. Note that, if $b>1$, $j_{*}c^{+}+j_{*}c^{-}=0$ in $H^{\operatorname{BM}}_{1}(G\setminus\{e\})$. Summarizing:
@@
{tikzpicture}
Figure 1. In the case $(a,b)=(1,1)$, the homology classes $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}Y$ in red and $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}W$ in blue. The pushforward of the Borel Moore homology classes $c^{+},c^{-}\in H_{1}^{\operatorname{BM}}(\mathbb{R}\setminus\{0\})$.
Remark 3.1.
Consider a central arrangement $\mathcal{A}$, the linear relations among the elements $\chi_{i}\in\Lambda\simeq\mathbb{Z}^{r}$ define a representable matroid. For any subvariety $H_{i}=\ker(\chi_{i})=\ker(-\chi_{i})$ we choose one of the two descriptions. These choices determine an oriented matroid.
We will denote the circuits of the matroid by $C$ and the oriented circuits by $C=C^{+}\sqcup C^{-}$.
Now we consider $\mathcal{A}$ as a real arrangement, for any $i\in E$ let $w_{i}^{+}=\chi_{i}^{*}(w^{+})$ and $w_{i}^{-}=\chi_{i}^{*}(w^{-})$ be the classes in $H^{0}(M^{\mathbb{R}}(\mathcal{A}))$ corresponding to the inverse image of $\mathbb{R}_{+}$ and $\mathbb{R}_{-}$ through $\chi_{i}\colon\mathbb{R}^{r}\rightarrow\mathbb{R}$.
@@
Let $\mathcal{A}$ be a central real arrangement, the ring $H^{0}(M^{\mathbb{R}}(\mathcal{A}))$ is generated by the classes $w_{i}^{+},w_{i}^{-}$ with $i\in E$ and subject to the following relations:
@@
Remark 3.3.
From the first and second relation it follows that $(w_{i}^{\pm})^{2}=w_{i}^{\pm}$. We consider a central circuit $C$ so that also the opposite $-C$ is a circuit. Then, using the first and the third relation for $C$, we get
@@
By summing the equations above with an opportune sign, we get the following relations:
$\sum_{J\subsetneq C^{-}}(-1)^{|J|}w^{+}_{C^{+}\sqcup J}-(-1)^{|C|}\sum_{J\subsetneq C^{+}}(-1)^{|J|}w^{+}_{J\sqcup C^{-}}=0$
$\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}w^{+}_{C\setminus K}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}w^{+}_{C\setminus K}=0$ (8)
From this identities, we will obtain a relation in $H^{*}(M^{a,b}(\mathcal{A}))$ by applying the map $i_{*}$.
@@
Let $A,B$ be disjount subsets of $E$. We denote by $\ell(A,B)$ the sign of the permutation taking the $A\sqcup B$ as a ordered subset of $E$ of $\mathcal{A}$ to the concatenation of $A$ and $B$.
@@
Let $B\subseteq E$ a basis of the matroid, we denote by $\operatorname{sgn}(B)$ the sign of the determinant of the matrix whose columns represent the elements $\chi_{b}$ (in a positive basis of $\Lambda$) with the order induced by $E$.
Remark 3.6.
Let $\mathcal{B}\subset\mathcal{A}$ be a sub-arrangement and $j\colon M^{a,b}_{\mathcal{A}}\hookrightarrow M^{a,b}_{\mathcal{B}}$ be the inclusion of the complements. From the definitions follow that $j^{*}\eta^{\mathcal{B}}_{C,D}=\eta_{C,D}^{\mathcal{A}}$. So in the following we will not specify the ambient space.
@@
Let $\mathcal{A}$ be a totally unimodular central arrangement of rank $n$. Consider the closed immersion $i\colon M^{\mathbb{R}}(\mathcal{A})\rightarrow M^{a,b}(\mathcal{A})$ and its pushforward in cohomology $i_{*}\colon H^{0}(M^{\mathbb{R}}(\mathcal{A}))\rightarrow H^{n(a+b-1)}(M^{a,b}(\mathcal{A}))$. Let $I\subseteq E$ be an independent set and $B\subseteq E$ a basis containing $I$:
@@
  • If $b>1$, then $i_{*}(w_{I}^{+})=\begin{cases}0&\text{ if }\operatorname{rk}(I)\neq n,\\ \operatorname{sgn}(I)^{d}\omega_{I}&\text{ otherwise.}\end{cases}$
@@
Let $\mathcal{B}\subset\mathcal{A}$ be the sub-arrangement $\{H_{b}\mid b\in B\}$. Consider the inclusions $j^{\mathbb{R}}\colon M^{\mathbb{R}}(\mathcal{A})\hookrightarrow M^{\mathbb{R}}(\mathcal{B})$ and $j^{a,b}\colon M^{a,b}(\mathcal{A})\hookrightarrow M^{a,b}(\mathcal{B})$. Recall that $d=a+b-1$ is the codimension of $\mathbb{R}$ in $G$. The following diagram commutes by the naturality of pushforward in cohomology (property (3)):

### v2

@@ Abstract.
We give a Orlik-Solomon type presentation for the cohomology ring of arrangements in a non-compact abelian Lie group. The new insight consists in comparing arrangements in different abelian groups. Our work is based on the Varchenko-Gel’fand ring for real hyperplane arrangements and from that we deduce the cohomology rings of all other abelian arrangements. As by-product, we obtain a new proof of the Orlik-Solomon relations and De Concini-Procesi ones.
1991 Mathematics Subject Classification:
14N20 (primary), 52C35, 55R80 (secondary)
@@ 1. Introduction
The topology of hyperplane arrangements is a classical subject whose study goes back to the sixties. It is originally motivated by the relations with braid groups [Del72] and configuration spaces [CLM76]. The study of arrangements was also motivated by combinatorics, in particular matroid theory, and by singularity theory.
A first generalization of hyperplane arrangements – now called toric arrangements – were introduced in [Loo93]. In the 2000s, De Concini and Procesi [DCP05] started a systematic study of toric arrangements motivated by the application to knapsack problem. Finally, abelian arrangements were introduced in 2016 by Bibby [Bib16].
An important focus on this subject is the description of the cohomology ring of the arrangement complement. Arnold identified relations between differential forms in the cohomology ring of braid hyperplane arrangements [Arn69] and Brieskorn described the cohomology ring of arrangements of Coxeter type [Bri72]. A milestone is the famous result by Orlik and Solomon [OS80]: they presented the cohomology ring of the complement of complex hyperplane arrangements by generators and relations. Later, Varchenko and Gel’fand studied the cohomology ring of real hyperplane arrangements using Heaviside functions as generators [VG87]. They observed an analogy between the real and the complex case:
@@
Proudfoot [Pro06] relates the VG-algebra and the OS-algebra with $\mathbb{F}_{2}$-coefficients overcoming the non-commutativity of the OS algebra. Moseley [Mos17] shows an isomorphism between the VG-algebra and the cohomology of (subspace) arrangements in $(\mathbb{R}^{3})^{n}$. The Varchenko-Gel’fand presentation was generalized to oriented matroids and conditional oriented matroids [GR89, Cor02, DB23, DBPW22].
A Orlik-Solomon type presentation for toric arrangements was obtained by De Concini and Procesi [DCP05] in the unimodular case and later extended to all toric arrangements [CD17, CDD+20]. The cohomology of the complement can be also computed, via the Morgan model, from the cohomology of a wonderful compactification and their strata. This approach, for toric arrangements, was developed in a sequence of articles [DCG18, DCG19, DCGP20, DCG21, MP22, GPS24].
@@
We obtain a new proof of the Orlik-Solomon result [OS80] for complexified hyperplane arrangements and of the De Concini-Procesi presentation [DCP05] for unimodular toric arrangements. We also obtain a new proof of the presentation of the rational cohomology ring for general toric arrangements, that firstly appeared in [CDD+20]. In all other cases our result is original.
We start by setting the notations in Section 2. In Section 3, we explicitly describe the pushforward along the map $i\colon M^{\mathbb{R}}(\mathcal{A})\to M^{a,b}(\mathcal{A})$ induced by the inclusion $\mathbb{R}\hookrightarrow\mathbb{R}^{b}\times(S^{1})^{a}$. We manipulate the Varchenko-Gel’fand relations to obtain relations in the cohomology ring of central and unimodular arrangements (see Lemma 3.6). Section 4 extends the results of [LTY21] to the case of non-central arrangements (see Theorem 4.1). Then, in Section 5 we use the technique of separating coverings developed in [CDD+20] to obtain the main theorem. In Section 6 we consider toric arrangements and we compare the presentation in [CDD+20] with our own; we also show how to derive the former from the latter. Finally, in Section 7 we apply our theorem to the arrangements of type $A_{n}$ improving a result of [CT78] on configuration spaces in $\mathbb{R}^{b}\times(S^{1})^{a}$.
@@ 2. Definitions and notations
Let $\Lambda\cong\mathbb{Z}^{r}$ be a lattice with the choice of an orientation, so that $\operatorname{Hom}(\Lambda,G)\simeq G^{r}$ has a natural orientation. A primitive character $\chi\in\Lambda$ corresponds to a group morphism $\chi\colon\operatorname{Hom}(\Lambda,G)\rightarrow G$. Let $E$ be a finite set with a total ordering. An abelian arrangement is a finite collection of connected subvarieties in $\operatorname{Hom}(\Lambda,G)$ defined by a finite list of primitive characters $\chi_{i}$ and elements $g_{i}\in G$ indexed by $E$:
@@
This definition, specialized to the case $G=\mathbb{C}$, corresponds to hyperplane arrangements with the property that each hyperplane has a normal vector with rational coordinates.
The poset of layers $\mathcal{L}(\mathcal{A})$ (or intersection poset) of $\mathcal{A}$ is the set of all connected components of intersections of elements in $\mathcal{A}$, ordered by reverse inclusion. An arrangement is called central if the intersection of all the subvarieties is non-empty, and it is called unimodular whenever all the possible intersections of elements in $\mathcal{A}$ is either connected or empty.
@@
$M(\mathcal{A})=\operatorname{Hom}(\Lambda,G)\setminus\bigcup_{i\in E}H_{i}.$
@@
Central abelian arrangements are related to matroids, oriented matroids, and arithmetic matroids, we briefly recall definitions and terminologies. For further details we refer to the standard textbooks [Oxl11, BLVS+99, And25] for oriented matroids, to the articles [DM13, BM14] for arithmetic matroids and to [Pag20, PP21] for the interaction between oriented and arithmetic matroids.
Consider a central abelian arrangement $\mathcal{A}=\{H_{i}=\chi_{i}^{-1}(e)\}_{i\in E}$, the collection of $\chi_{i}\in\Lambda$ define a realizable, oriented and arithmetic matroid in the following way. A matroid $(E,\operatorname{rk})$ is a pair of a groundset $E$ and a rank function $\operatorname{rk}\colon 2^{E}\to\mathbb{N}$ satisfying the three properties:
  1. (R1)
    $\operatorname{rk}(A)\leq\lvert A\rvert$ for all $A\subseteq E$,
  2. (R2)
    (increasing) $\operatorname{rk}(A)\leq\operatorname{rk}(B)$ for all $A\subseteq B\subseteq E$,
  3. (R3)
    (submodular) $\operatorname{rk}(A)+\operatorname{rk}(B)\geq\operatorname{rk}(A\cap B)+\operatorname{rk}(A\cup B)$ for all $A,B\subseteq E$.
For any central arrangement $\mathcal{A}$, consider the rank function $\operatorname{rk}_{\mathcal{A}}(A)=\operatorname{rk}_{\mathbb{Z}}\langle\chi_{i}\mid i\in A\rangle$ for any $A\subseteq E$; the pair $(E,\operatorname{rk}_{\mathcal{A}})$ is a matroid. The real codimension of the intersection $\cap_{i\in A}H_{i}$ is equal to $(a+b)\operatorname{rk}_{\mathcal{A}}(A)$. A subset $A$ is independent if $\operatorname{rk}(A)=\lvert A\rvert$, dependent if $\operatorname{rk}(A)<|A|$, and a basis if $\operatorname{rk}(A)=\lvert A\rvert=\operatorname{rk}(E)$. A subset $B$ is dependent on $A\subseteq E$ if $\operatorname{rk}(A\cup B)=\operatorname{rk}(A)$. A set $C\subseteq E$ is a circuit if it is a minimally dependent set, i.e. $\operatorname{rk}(C)=\lvert C\rvert-1$ and for all $A\subsetneq C$ $\operatorname{rk}(A)=\lvert A\rvert$. A subset $A\subseteq\mathcal{A}$ has nullity 111 if $\operatorname{rk}(A)=\lvert A\rvert-1$ and, in this case, we denote by $C(A)$ the unique circuit contained in $A$.
An arithmetic matroid is a triple $(E,\operatorname{rk},m)$ where $(E,\operatorname{rk})$ is a matroid and the multiplicity function $m\colon 2^{E}\to\mathbb{N}_{>0}$ satisfies:
  1. (AM1)
    $m(A\cup\{i\})\mid m(A)$ for all $A\subseteq E$ and all $i\in E$ dependent on $A$,
  2. (AM2)
    $m(A)\mid m(A\cup\{i\})$ for all $A\subseteq E$ and all $i\in E$ independent on $A$,
  3. (AM3)
    for all subsets $A\subseteq B\subseteq E$ such that $B$ is a disjoint union $B=A\sqcup F\sqcup T$ satisfying $\operatorname{rk}(C)=\operatorname{rk}(A)+\lvert C\cap F\rvert$ for all $A\subseteq C\subseteq B$, then
    $m(A)m(B)=m(A\cup F)m(A\cup T),$
  4. (AM4)
    for all $A\subseteq B\subseteq E$ with $\operatorname{rk}(A)=\operatorname{rk}(B)$, then
    $\sum_{A\subseteq C\subseteq B}(-1)^{\lvert C\setminus A\rvert}m(C)\geq 0,$
  5. (AM5)
    for all $A\subseteq B\subseteq E$ with $\lvert A\rvert+\operatorname{rk}(E\setminus A)=\lvert B\rvert+\operatorname{rk}(E\setminus B)$, then
    $\sum_{A\subseteq C\subseteq B}(-1)^{\lvert C\setminus A\rvert}m(E\setminus C)\geq 0.$
The function $m_{\mathcal{A}}\colon 2^{E}\to\mathbb{N}_{>0}$ defined by $m_{\mathcal{A}}(A)=\#\operatorname{tor}\Lambda/\langle\chi_{i}\mid i\in A\rangle$, i.e. the cardinality of the torsion of the quotient $\Lambda/\langle\chi_{i}\mid i\in A\rangle$ endows the matroid $(E,\operatorname{rk}_{\mathcal{A}})$ of an arithmetic matroid structure. Since the arrangement $\mathcal{A}$ is central, the number of connected components of the intersection $\cap_{i\in A}H_{i}$ is equal to $m_{\mathcal{A}}(A)^{a}$. In particular, for $a=0$ all intersections are connected. A subset $A\subseteq E$ is unimodular if $m(A)=1$.
The above definitions do not change if we change the description of some $H_{i}$ replacing $\chi_{i}$ with $-\chi_{i}$. Now we fix for each subvariety $H_{i}$ a choice of a character $\chi_{i}$, the construction of the oriented matroid will depend on these choices.
A signed subset $(C^{+},C^{-})$ of $E$ is a pair of disjoint subset of $E$ and we denote its support by $C=C^{+}\sqcup C^{-}$. An oriented matroid is a pair $(E,\mathcal{C})$ of a groundset $E$ and a collection of signed subsets of $E$ such that:
  1. (C0)
    $(\emptyset,\emptyset)\not\in\mathcal{C}$,
  2. (C1)
    (symmetry) if $(C^{+},C^{-})\in\mathcal{C}$, then $(C^{-},C^{+})\in\mathcal{C}$,
  3. (C2)
    (incomparability) for all $(C^{+},C^{-}),(D^{+},D^{-})\in\mathcal{C}$ such that $C\subseteq D$ then $(C^{+},C^{-})=(D^{+},D^{-})$ or $(C^{+},C^{-})=(D^{-},D^{+})$,
  4. (C3)
    (circuit elimination) for all $(C^{+},C^{-}),(D^{+},D^{-})$ such that $(C^{+},C^{-})\neq(D^{-},D^{+})$ and for all $i\in C^{+}\cap D^{-}$ there exists $(B^{+},B^{-})\in\mathcal{C}$ such that $B^{+}\subseteq(C^{+}\cup D^{+})\setminus\{i\}$ and $B^{-}\subseteq(C^{-}\cup D^{-})\setminus\{i\}$.
We call oriented circuits the signed subsets of an oriented matroid. The set $\{C\}_{(C^{+},C^{-})\in\mathcal{C}}$ is the set of circuits of a matroid and such a matroid is unique.
Consider, for each minimal linear relation $\sum_{i\in C}n_{i}\chi_{i}=0$ among characters of $\mathcal{A}$, the sets $C^{+}=\{i\in E\mid n_{i}>0\}$ and the set $C^{-}=\{i\in E\mid n_{i}<0\}$. The collection $\mathcal{C}_{\mathcal{A}}$ of all such pairs $(C^{+},C^{-})$ defines an oriented matroid. Notice that for each signed circuit $(C^{+},C^{-})$ the linear relation is unique up to scalar multiplication, so also the pair $(C^{-},C^{+})$ is a signed circuit and we call it the opposite circuit of $(C^{+},C^{-})$. For the sake of notation we will denote an oriented circuit by $C=C^{+}\sqcup C^{-}$.
An alternative definition of oriented matroids is given by basis orientations (also called chirotopes): let, for each basis $B\subseteq E$, $\operatorname{sgn}(B)$ the sign of the basis $(b_{1},b_{2},\dots,b_{r})$ of $\Lambda\otimes_{\mathbb{Z}}\mathbb{Q}$, where we take the element of $B=\{b_{1},\dots,b_{r}\}$ in the order prescribed by the ordered set $E$. The collection of $\operatorname{sgn}(B)$ for all basis $B$ of the matroid determines the oriented matroid. Changing the orientation of $\Lambda$ reverses all signs $\operatorname{sgn}(B)$, but does not affect the oriented matroid.
The situation for non-central arrangements is different. Let $\mathcal{A}=\{H_{i}=\chi_{i}^{-1}(g_{i})\}$ be an abelian arrangement, the collection of characters $\chi_{i}$ for $i\in E$ defines an oriented arithmetic matroid as in the central case. For $A\subseteq E$ if the intersection $\cap_{i\in I}H_{i}$ is non-empty, then it has codimension $(a+b)\lvert A\rvert$ and $m(A)^{a}$ connected components. If $A$ is an independent set, the intersection is certainly non-empty. Otherwise, it depends on the elements $g_{i}$ for $i\in A$. We say that a dependent set is central if the corresponding intersection is non-empty.
@@
In order to define certain cohomological classes in $H^{a+b-1}(G\setminus\{e\})$ we consider the immersion of a small sphere $i\colon S^{a+b-1}\hookrightarrow G\setminus e$ centered in $e$. Let $W$ be the top dimensional class generating $H_{a+b-1}(S^{a+b-1})$. We denote by $\omega$ the class in $H^{a+b-1}(G\setminus\{e\})$ dual to $i_{*}(W)$ with the orientation given by the outer normal first rule.
@@
$i_{x}\colon\quad(S^{1})^{a}$ $\hookrightarrow G\setminus\{e\}$
@@
Up to homotopy, $i_{x}$ only depends on the connected component of $\mathbb{R}^{b}\setminus\{0\}$ the point $x$ belongs to. In the case $b=1$ we choose $x<0$, otherwise the choice of the point does not matter. Denote by $\psi^{j}$ the class in $H^{1}(G\setminus\{e\})$ dual to $(i_{x})_{*}(Y^{j})$ and by $\psi$ the class in $H^{a}(G\setminus\{e\})$ dual to $(i_{x})_{*}(Y)$.
For any $i\in E$, the map $\chi_{i}$ induces $(\chi_{i}-g_{i})|_{M^{a,b}(\mathcal{A})}\colon M^{a,b}(\mathcal{A})\rightarrow G\setminus\{e\}$. So we set
@@
We have $\omega_{i}\psi_{i}^{j}=0$ for any $i\in E$ and any $j\leq a$ because $\omega\psi^{j}=0$ in $H^{*}(G;\mathbb{Z})$. In particular, $\omega_{i}z=0$ for any $z\in\ker\left(H^{*}(G^{r};\mathbb{Z})\to H^{*}(H_{i};\mathbb{Z})\right)$. Note that, when $b=1$, the classes $\psi_{i}$’s have degree $a=a+b-1$, the same degree of the elements $\omega_{i}$’s.
@@ 3. Central and unimodular arrangements
Consider a proper map $f\colon N\to M$ between oriented manifolds. Such a map induces a pushforward map in cohomology $f_{*}$ obtained from the pushforward in Borel-Moore homology (for a general reference see [HR96]) by composing with the Poincaré duality isomorphism.
[…]
@@
We consider the pushforward in cohomology $i_{*}\colon H^{0}(\mathbb{R}\setminus\{0\})\rightarrow H^{a+b-1}(G\setminus\{e\})$ induced by the closed inclusion $i\colon\mathbb{R}\setminus\{0\}\rightarrow G\setminus\{e\}$, $i(x)=(x,0,\cdots,0,1,\cdots,1)$. Let $w^{+}$ and $w^{-}$ be the two standard generators of $H^{0}(\mathbb{R}\setminus\{0\})$. Poincaré Duality maps $w^{+}$ and $w^{-}$ to the classes in $H^{\operatorname{BM}}_{1}(\mathbb{R}\setminus\{0\})$ represented by the infinite chains $c^{+}$ and $c^{-}$, corresponding respectively to the positive and negative semi-axes. By checking intersection numbers, we obtain $\operatorname{PD}(\omega)=i_{*}c^{+}$ and, when $b=1$, $\operatorname{PD}(\psi)=i_{*}c^{+}+i_{*}c^{-}=i_{*}1$. Note that, if $b>1$, $i_{*}c^{+}+i_{*}c^{-}=0$ in $H^{\operatorname{BM}}_{1}(G\setminus\{e\})$. Summarizing:
@@
$Y$$W$$i_{*}c^{-}$$i_{*}c^{+}$(×$(\mathbb{R}\times S^{1})\setminus(0,1)$
Figure 1. In the case $(a,b)=(1,1)$, the homology class $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Y$ is represented in red and $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}W$ in blue. The two arrows represent the support of Borel Moore homology classes $i_{*}c^{+},i_{*}c^{-}\in H_{1}^{\operatorname{BM}}(\mathbb{R}\setminus\{0\})$.
Now we consider $\mathcal{A}$ as a real hyperplane arrangement, for any $i\in E$ let $w_{i}^{+}=\chi_{i}^{*}(w^{+})$ and $w_{i}^{-}=\chi_{i}^{*}(w^{-})$ be the classes in $H^{0}(M^{\mathbb{R}}(\mathcal{A}))$ corresponding to the positive and negative half spaces.
@@
Let $\mathcal{A}$ be a central real hyperplane arrangement. The ring $H^{0}(M^{\mathbb{R}}(\mathcal{A}))$ is generated by the classes $w_{i}^{+},w_{i}^{-}$ with $i\in E$ and subject to the following relations:
@@
Remark 1.
From the first and second relations it follows that $(w_{i}^{\pm})^{2}=w_{i}^{\pm}$. Consider a central circuit $C$, using the first and the third relations for $C$, we get
@@
By summing the equations above with a suitable sign, we get the following relations:
$\sum_{J\subsetneq C^{-}}(-1)^{|J|}w^{+}_{C^{+}\sqcup J}-(-1)^{|C|}\sum_{J\subsetneq C^{+}}(-1)^{|J|}w^{+}_{J\sqcup C^{-}}=0,$
$\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}w^{+}_{C\setminus K}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}w^{+}_{C\setminus K}=0.$ (4)
From this identities, we will obtain a relation in $H^{*}(M^{a,b}(\mathcal{A}))$ by applying the pushforward map $i_{*}$.
@@
Let $A,B$ be disjoint subsets of $E$. We denote by $\ell(A,B)$ the sign of the permutation taking the $A\sqcup B$ as a ordered subset of $E$ of $\mathcal{A}$ to the concatenation of $A$ and $B$.
@@
Remark 2.
Let $\mathcal{B}\subset\mathcal{A}$ be a sub-arrangement and $j\colon M^{a,b}_{\mathcal{A}}\hookrightarrow M^{a,b}_{\mathcal{B}}$ be the inclusion of the complements. From the definitions it follows that that $j^{*}\eta^{\mathcal{B}}_{C,D}=\eta_{C,D}^{\mathcal{A}}$. So we will not specify the ambient space. Since $\omega_{A}z=0$ for any $z\in\ker\left(H^{*}(G^{r};\mathbb{Z})\to H^{*}(\cap_{i\in C}H_{i};\mathbb{Z})\right)$, we have $\eta_{C,D}=0$ if $D$ is dependent on $C$.
@@
Let $\mathcal{A}$ be a unimodular central arrangement of rank $r$. Consider the closed immersion $i\colon M^{\mathbb{R}}(\mathcal{A})\rightarrow M^{a,b}(\mathcal{A})$ and its pushforward in cohomology $i_{*}\colon H^{0}(M^{\mathbb{R}}(\mathcal{A}))\rightarrow H^{r(a+b-1)}(M^{a,b}(\mathcal{A}))$. Let $I\subseteq E$ be an independent set and $B\subseteq E$ a basis containing $I$:
@@
  • If $b>1$, then $i_{*}(w_{I}^{+})=\begin{cases}0&\text{ if }\operatorname{rk}(I)\neq r,\\ \operatorname{sgn}(I)^{d}\omega_{I}&\text{ otherwise.}\end{cases}$
@@
Let $\mathcal{B}\subset\mathcal{A}$ be the sub-arrangement $\{H_{b}\mid b\in B\}$. Consider the inclusions $j^{\mathbb{R}}\colon M^{\mathbb{R}}(\mathcal{A})\hookrightarrow M^{\mathbb{R}}(\mathcal{B})$ and $j\colon M^{a,b}(\mathcal{A})\hookrightarrow M^{a,b}(\mathcal{B})$. Recall that $d=a+b-1$ is the codimension of $\mathbb{R}$ in $G$. The following diagram commutes by the naturality of pushforward in cohomology (property (P3)):